a flywheel of gasoline engine is required to give up 300 J of kinetic energy while its angular velocity decreases from 600 rev/min to 540 rev/min. What moment of inertia is required?


Grade:

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear cathy

E  =1/2 Iw12  - 1/2 Iw22

  =I/2 [w12  - w22]

300 = I/2 [ w12  - w22]

 

w1 = 2*∏*600 /60 = 20∏ rad/sec

w2 = 2*∏*540 /60 = 18∏ rad/sec

 

put the value of w1 and w2 and calculate the value of moment of Inertia

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.